OpenStudy (anonymous):

What is the point-slope form of the equation of a line that passes through the point (6, −8) and has a slope of −2?

OpenStudy (texaschic101):

point slope form : y - y1 = m (x - x1) slope (m) = -2 (6,-8)...x1 = 6 and y1 = -8 now we sub y - (-8) = -2 (x - 6) y + 8 = -2 ( x - 6) <====point slope form